OpenStudy (anonymous):

Reduce -34/87 to lowest terms. a) -2/5 b) -1/3 c) -11/29 d) It is already reduced.

OpenStudy (anonymous):

Can u find GCF of 34 and 87

OpenStudy (anonymous):

I'm pretty sure It's already reduced :P

OpenStudy (anonymous):

The only factor of 87 is 3, and you can't divide 34 by 3 unless it's a decimal :P Might be wrong! Pretty sure though

OpenStudy (anonymous):

34 =2*17 87 = 3 * 29 Thus, the GCF of 34 and 87 is 1.

OpenStudy (anonymous):

So, it is already reduced

OpenStudy (anonymous):

@sarah.radke factor of 87 is 29 also not only 3

OpenStudy (anonymous):

@sauravshakya oh, I totally knew that lol :P